OpenStudy (anonymous):

Simplify the expression: 4x2 − 8y + 3x − 9x2 + 10y

OpenStudy (mathstudent55):

The terms with x^2 are like terms and can be combined/ The terms with y are like terms and can be combined. Then add the term with x

OpenStudy (mathstudent55):

4x2 − 8y + 3x − 9x2 + 10y = 4x^2 - 9x^2 + 10y - 8y + 3x = ?

OpenStudy (anonymous):

You have to combine like terms. So basically. 4x^2 - 9x^2 = -5x^2 and -8y +10y=2y You cannot combine x because there arent any like terms for x. So your final answer will be -5x^2 + 3x +2y

OpenStudy (anonymous):

4x^2 - 8y + 3x - 9x^2 + 10y

OpenStudy (anonymous):

thanks safira97

OpenStudy (anonymous):

4x^2 -8y+3x-9x^2+10y =4x^2-9x^2+3x-8y+10y =-5x^2+3x+2y